shift summary evening shift Dec 31,1999


Subject: shift summary evening shift Dec 31,1999
From: Vivian O'Dell (odell@fnal.gov)
Date: Wed Jan 12 2000 - 16:24:34 CST


 Shift summary for the last shift of 1999.

 On shift: Mike Arenton

 Summary: very little happening. Run 15349 is running at end of shift
with only calibration triggers. Earlier tried to turn on beam cosmic
triggers but got lots of errors, not fixed by usual KFREND's etc.

 DN changed CO2 to cylinder since dewar was empty.



This archive was generated by hypermail 2b27 : Wed Jan 12 2000 - 16:24:45 CST